A penny ,a nickel and a dime are tossed simultaneously. What is the probability that exactly 2 of the coins show heads when they land?
Answer by venugopalramana(3286)

A penny ,a nickel and a dime are tossed simultaneously. What is the probability that exactly 2 of the coins show heads when they land?
probabilty of head or tail coming up for each coin = 0.5
since 3 coins are different,the total number of possibilities =2*2*2=8
we need any 2 to come up heads and the other to come up tail.
selecting 2 for heads can be done in 3C2=3 ways.
probability =3/8
alternately we can find probability by
3 ways *p of head * p of head * p of tail = 3*(1/2)*(1/2)*(1/2)